21) Order for Tobramycin 50mg IV q8h. The female patient weighs 143 lbs and the physician states he wants 3mg/kg/day to avoidPlease explain parts a-f. I appreciate all of the help!
Answer #1

a) patient weight 143 lbs

weight in kg=143/2.2=71.5 kg

b)50 mg q8h then for per day =150mg

c) safe limit is 3mg/kg/day

for 71.5 mg odse will be=214.5 mg per day

d) yes it is safe order as the ordered 150mg perday is lesser than the limit i.e. 214.5 mg per day

e. available is 80mg/2ml=40mg/ml

required dose 50mg

then required ml=1.25 ml q 8h

f)pump rate in ml/min=600/40

=15ml/min
Answer #2

a) patient weight 143 lbs

weight in kg=143/2.2=71.5 kg

b)50 mg q8h then for per day =150mg

c) safe limit is 3mg/kg/day

for 71.5 mg odse will be=214.5 mg per day

d) yes it is safe order as the ordered 150mg perday is lesser than the limit i.e. 214.5 mg per day

e. available is 80mg/2ml=40mg/ml

required dose 50mg

then required ml=1.25 ml q 8h

f)pump rate in ml/min=600/40

=15ml/min